Required: 1. A partially completed schedule of the companys total and per unit costs over the relevant range of 63,000 to 103,000 units produced and sold annually is given below: Complete the schedule of the companys total and unit costs. (Round the variable cost and fixed cost to 2 decimal places.) Units Produced and Sol 63,000 83,000 103,000 Total costs: Variable costs $ 252,000 370,000 Fixed costs 370,000 370,000 $ 622,000 370,000370,000 Total costs Cost per unit Variable cost Fixed cost Total cost per unit 0.00 $ 0.00 S 0.00
Assume that the company produces and sells 93,000 units during the year at a selling price of $9.46 per unit. Prepare a contribution format income statement for the year Harris Company Contribution Format Income Statement
Answer #1
Units Produced
63,000 83,000 1,03,000
Total Costs:
    Variable Costs $ 252,000 $ 332,000 $ 412,000
    Fixed Costs $ 370,000 $ 370,000 $ 370,000
Total Costs: $ 622,000 $ 702,000 $ 782,000
Cost per Unit:
Variable Cost $ 4.00 $ 4.00 $ 4.00
Fixed Cost $ 5.87 $ 4.46 $ 3.59
Total Cost per unit $ 9.87 $ 8.46 $ 7.59
Notes to Schedule:
Variable Cost per unit = Total variable Costs ÷ Units produce
V.C per unit = $252,000 ÷ 63,000 = $ 4 per unit at each level of output
Fixed Cost per unit= Total Fixed Cost ÷ Level of output
F.C per unit at 63,000 of output = $370,000 ÷ 63,000 = $ 5.87 per unit
F.C per unit at 83,000 of output = $370,000 ÷ 83,000 = $ 4.46 per unit
F.C per unit at 103,000 of output = $370,000 ÷ 103,000 = $ 3.59 per unit
Harris Company
Contribution Format Income Statement
Sales Revenue (93,000 units @ $9.46 per unit) $ 879,780
Less: Variable Costs: (93,000 units @ $4 per unit) $ 372,000
           Contribution Margin $ 507,780
Less: Fixed Costs $ 370,000
Operating Income/Profit $ 137,780
Fixed Cost per unit= Total Fixed Cost ÷ Level of output
At 93,000 level of Output= $370,000 ÷ 93,000 = $3.98 per unit